OpenStudy (anonymous):

A rectangle has an area of 25 square centimeters and a length of 10 centimeters. What is the width of the rectangle? A. 2.5 centimeters B. 35 centimeters C. 7.5 centimeters D. 15 centimeters

OpenStudy (anonymous):

@sissyedgar @ribhu

OpenStudy (trojanpoem):

Area= length * width 25 = 10 * width width = 2.5 centimeters

OpenStudy (ribhu):

25=10*x; x=25/10=2.5

OpenStudy (sissyedgar):

I agree wit 2.5:P

OpenStudy (anonymous):

Thanks all of you